# Dual comb–enhanced microwave clock synchronization over commercial fiber

### v1

@@ Abstract
The large-scale clock network is the key ingredient to obtain high precision in many scenarios, from fundamental research to cutting-edge applications. However, time synchronization among microwave clocks has become an upper limit of the time-frequency network because of the inherent constrained precision in optical pulse measurements. Here, we demonstrate a femtosecond-level time synchronization of microwave clocks through a commercial link of 205.86 km via dual comb–enhanced optical two-way time transfer, which achieves a 6.23-fs residual time deviation between synchronized timescales at 1 s and an instability below at 10,000 s. Further, the high-precision time synchronization of microwave clocks significantly enhances the probe ability of subtle reciprocity changes of fiber to the sub-picosecond level. This work provides a path toward secure fiber time-frequency networks to support future microwave clock–based precise timing and sensing systems.
@@ I Introduction
High-precision clock networks offer the fundamental technical foundation for numerous scenarios, including positioning, navigation, timing [1, 2, 3], sensitive probes [4, 5, 6], and fundamental physics testing [7, 8, 9, 10]. The enormous benefits are unleashed through the full interconnection between any two nodes in a hybrid network of optical clocks and microwave clocks. In particular, the time-frequency transfer over both fiber and free-space optical links has seen tremendous progress [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. The ultra-high precision frequency comparison has been achieved using well-established optical frequency transfer techniques by directly disseminating stable frequency signals [14, 15, 16, 25, 26], whereas the timescale comparison requires an exchange of a series of labeled optical pulses generated by frequency standards [11, 12, 13, 17, 18, 19, 20, 21, 22, 23, 24]. Since the optical two-way time transfer (OTWTT) method uses the reciprocity of bi-directional light travel to eliminate most of the common-mode noises and the link drift [31, 32], the measurement of optical pulses has become the key to improving the precision for time transfer.
@@
In the time synchronization of two microwave clocks, the electrical measurement has not been replaced by the optical approach because a phase coherence of independent optical frequency combs (called combs for simplicity) at both local and remote sites, which is a necessary requirement for performing stable and low-jitter interference, is challenging to establish. Specifically, microwave-to-microwave synchronization is difficult because both reference and synchronized clocks have greater phase noise than optical clocks and the pulse coherence is reduced by propagation-induced pulse impairments.
To overcome the limitations imposed by measurement devices on the time synchronization of microwave clocks, here, we propose a tight microwave-clock time synchronization scheme based on the dual comb–enhanced OTWTT and provide a demonstration over 205.86 km commercial fiber. We achieved a 6.23-fs residual time deviation between synchronized timescales at 1 s and an instability below at 10,000 s due to a series of breakthroughs. The intensity-encoded comb pulses were measured using a Kalman filter–assisted LOS, where the centroids of reference and target interferograms at each site served as the start and stop times of the time interval, to reduce the impact of the intensity noise and pulse distortion on the peak measurements. We used the phase time of the pulse envelope instead of the carrier phase for time synchronization [20] to remove the limit of the carrier envelope offset (CEO) frequency, since the stable carrier phase is vulnerable to fluctuations in independent sources and fiber transmission. Dispersion compensation, fully automatic polarization control, and real-time feedback are essential to guarantee the mutual optical coherence.
@@
The improvement in microwave clock synchronization capabilities has greatly expanded the application range of the entire high-precision clock networks.
II Results
Experimental layout. The experiment was performed in Wuhan commercial link. As shown in Fig. 1 (a), two sites A and B are adjacent to each other and located at Wuhan laboratory with independent rubidium clocks to easily quantify the residual transfer noise. The master site (site A) and the synchronized site (site B) were connected through two geographically folded optical fibers linked in Liuxinji (round-trip node), and the total link length was 205.86 km (the total link loss was 57.18 dB). The long-haul fiber transmission necessitated the installation of two optical repeater stations at Wuhan North, each of which contained an erbium-doped fiber amplifier (EDFA) and a dispersion compensating fiber (DCF) along the terrestrial link to recover the signal impairments caused by loss and dispersion. OTWTT was achieved by using the dense wavelength division multiplexing (DWDM) technology over the shared fiber link.
@@ II Results
Dual comb–based LOS measurement. At each site, we used a dual comb–based LOS measurement to replace the TIC in the measurement of pulses generated by microwave clocks. As shown in Fig. 1 (b), at each site, we first created a local timescale by phase-locking the repetition frequency of the 100-MHz comb to the local microwave clock, which served as the clock comb. Then, a 100.001-MHz comb, which was also locked to the local clock, output optical pulses with a repetition frequency offset by kHz relative to the clock comb, which served as the local oscillator (LO) comb to sample clock combs for performing LOS, as shown in Fig. 1 (c). Phase locking was achieved with a locking-induced noise of 45 fs (time deviation at 1 s) through an asynchronous locking system (see Methods).
@@
Both combs had center frequencies of 1,560 nm and a 3-dB spectrum bandwidth of approximately 15 nm. The spectral effectively covered the commonly used C-band range in fiber optic communication, which is necessary for two-way transmission. Here, we directly applied the international-telecommunication-union-standard (ITU) DWDM module to the output of each comb to generate band-limited pulses with different spectra. We filtered the LO comb to generate two outputs with a bandwidth of 0.8 nm (100 GHz), which were centered at 1,551.72 nm (ITU-C32) and 1,550.12 nm (ITU-C34). We filtered the clock comb to generate an output centered at ITU-C32 (for Site A) or ITU-C34 (for Site B). Since directly using the LOS for synchronization will introduce the ambiguity of ns, which causes cycle slips, we provided an unambiguous synchronization method by encoding time markers on the optical pulse using Mach-Zehnder modulator (MZM) (Methods). For each site, the output after the intensity modulator was separated into two parts. One part was sent to the link for transmission with the output power of 1 mW; the other part was to generate the local reference interferogram.
@@
At each site, two cross-correlations were generated. The reference interferogram was generated by the cross-correlation between the LO comb and the local clock comb. The target interferogram was generated by the cross-correlation between the LO comb and the incoming clock comb from the remote site, which correspond to the “time-stretched” start and stop signals of the measuring time intervals, respectively. The cross-correlations were generated with the equivalent sampling rate THz in the optical frequency domain and the updating rate of ms in the laboratory timescale. The equivalent timescale stretching parameter was obtained as . Then, we implemented an analog-to-digital converter (ADC) with a 400-MHz sampling rate and a 10-bit resolution to digitize the train of interferograms. The data were subsequently transmitted to a field programmable gate array (FPGA) controller and synchronized with the local clock for data processing.
@@
Timing discrimination. The fundamental task of OTWTT is to recover timing information from exchanged optical pulses. Because signal distortion and noise accumulation were caused by the long-haul fiber transmission, high-precision timing discrimination was challenging.
Instead of the commonly used heterodyne-voltage-to-time scaling [12], peak location [13] or carrier phase [20] to define the arrival time in free-space scenarios, we used the centroid of the interferogram to discriminate the temporal position of the optical pulses, which is defined by
for the -th labeled interferogram. Here, is the envelope extracted by the Hilbert transformation, since it is not sensitive to the pulse shape and can smooth out the intensity noise superimposed on optical pulses. Then, the time difference was obtained by comparing the centroid position of two interferograms.
@@
Kalman filter–assisted LOS measurement. In contrast to the signal fading from turbulence in the free-space scenario [12, 19], the decrease in visibility of the target interferogram mainly attributes to the polarization disturbance in fiber, which is eliminated by the in-line electronic polarization controller. Therefore, the hold-over behavior of the Kalman filter was not considered in our experiments. In this paper, we focused on the Kalman filter to limit the amount of measurement noise [41, 42]. We assumed two zero-mean Gaussian random processes in the prior noise model of Kalman filter to represent the random-walk phase noise and random-walk frequency noise for real-time data processing [22].
Fig. 2 (c) shows noise power spectral densities (PSDs) of the measurement process. The effective bandwidth of the Kalman filter was 0.01 Hz, and the PSD was reduced by approximately three orders of magnitude reduction at 0.5 Hz to 1.2 , which is close to the noise floor of the optical clock–based LOS measurement [12, 13, 22].
In addition, Figs. 2 (d) and (e) show that the Kalman filter–assisted LOS optimizes the measurement instability for a short averaging time. The shaded region shows the noise suppression below 100 s. In Fig. 2 (d), the time deviation of noise floor was optimized to 3.7 0.1 fs for 1 s of averaging time, which can support the measurement of femtosecond-level fiber changes. In Fig. 2 (e), the modified Allan deviation (MDEV) changed the slope of the noise floor from to after the Kalman filtering ( is the averaging time in seconds). Thus, the dominate white phase measurement noise was suppressed to the lower flicker phase noise region at short timescales (below 100 s), which determined the noise limitation of the time offset measurement.
In-loop time offset calculation. After the precision of the measurement part had been quantified, the real-time timing discrimination of both sites and calculation of the in-loop time offset were performed in the two FPGA controllers. First, at both sites, the respective time differences under laboratory timescale and were obtained by locally comparing the centroids of the labeled reference and target interferograms using the time difference calculation algorithm. The effective (actual) time differences at both sites are and . Then, the local generated time difference at site A was transmitted to site B using a modulated laser centered at 1,550.92 nm (ITU-C33) over the DWDM communication link. Combining with the local calculated time information , the in-loop time offset was calculated in the FPGA controller of Site B.
@@
Free-running link. First, we used our system to investigate the one-way free-running link. Specifically, we set both sites that shared a common master clock to test the open-loop link noise from A to B. After passing through the link, the incoming clock comb accumulated link effects, which resulted in the pulse impairments. One impairment is dispersion, which caused pulse broadening and distortion. The results show both second-order dispersion, which caused pulse broadening, and third-order dispersion, which resulted in pulse asymmetry and distortion. Additionally, the link noise and amplified spontaneous emission (ASE) noise introduced by EDFAs inevitably introduced intensity fluctuations in the optical pulses. Both dispersion-induced noise and ASE-induced intensity fluctuations decreased the pulse coherence, which ultimately affected the measurement jitter of the optical cross-correlation.
@@
Synchronized link. After fully characterizing the 205.86-km fiber link, by performing an OTWTT to cancel the link’s influence and closing the synchronized feedback loop, we obtained the performance of the synchronized link, which was continuously acquired for over 24 hours. As shown in Fig. 1 (d), the time offset calculation, which acted as the error signal, was fed into the digital feedback system that contained a proportional integral loop filter and a direct digital synthesizer (DDS). It also had a 30-Hz synchronization bandwidth to force the minimized residual timing jitter of the clock output to synchronize two microwave clocks.
@@
The residual MDEV between remote parties show the performance of the time-frequency transfer in Fig. 3 (b). The residual MDEV began at at 1 s and reached below at 10,000 s. The residual MDEV was below the state-of-the-art microwave clocks [44, 45, 46, 47] after 1 s of averaging time and more than two orders of magnitude better after the long averaging time. The slope was below 1,000 s, which indicates that the dominant noise at short timescales is flicker phase noise, and the residual MDEV follows power laws after 1,000 s averaging time. All of the results are consistent with the time deviation.
We also compared the stability result of in our experiment with the free-space scenarios that synchronized the optical clocks. The stability was approximately one order of magnitude less than that of Ref [13], given by a few . This lower starting point was mainly due to the higher noise floor of the LOS measurement and ASE noise of the in-line EDFAs, which relatively weakened the coherence of optical pulses and caused a lager timing jitter in the microwave-clock-synchronization scenarios than in the optical oscillator–based scenarios. Moreover, compare to the result of in Ref. [12], the trend of comes from the flicker phase noise caused by the technical noise of devices, so the delay-unsuppressed white-phase noise especially with low frequency [48, 49] could not be detected. Thus, we should optimize the microwave clock–based LOS measurement. Nevertheless, this result, which is close to the optical clock time synchronization, greatly closes the gap of microwave-microwave time synchronization in the entire time-frequency network.
Probing fiber delay non-reciprocity. Similar to the probe of non-reciprocity from multi-path effects [12] or motion platforms [19] in free space, many impacts can affect the non-reciprocity in fiber-based OTWTT, which affects the true clock-offset estimation results. A particular example of the recently extensively concerned potential asymmetric delay attacks [37, 50, 39]. The improvement of the clock-offset measurement enables us to sense the subtle delay non-reciprocity of the fiber link and improves the security of long-haul fiber transmission.

### v2

# Dual-comb-enhanced microwave clock synchronization over commercial fiber

@@ Abstract
The large-scale clock network is the key ingredient to obtain high precision in many scenarios, from fundamental research to cutting-edge applications. The advantage of the time synchronization among microwave clocks is their cost, size, and accessibility. Here, we demonstrate a femtosecond-level time synchronization of microwave clocks through a commercial link of 205.86 km via dual-comb-enhanced optical two-way time transfer, which achieves a 6.23-fs residual time deviation between synchronized timescales at 1 s and an instability below at 10,000 s. Further, the high-precision time synchronization of microwave clocks significantly enhances the probe ability of subtle reciprocity changes of fiber to the sub-picosecond level. This work provides a path toward secure fiber time-frequency networks to support future microwave-clock-based precise timing and sensing systems.
@@ I Introduction
High-precision clock networks offer the fundamental technical foundation for numerous scenarios, including positioning, navigation, timing [1, 2, 3], sensitive probes of environmental fluctuations [4, 5, 6], and fundamental physics testing [7, 8, 9, 10]. The enormous benefits are unleashed through the full interconnection between any two nodes in a hybrid network of optical clocks and microwave clocks. In particular, the time-frequency transfer over both fiber and free-space optical links has seen tremendous progress [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. The ultra-high precision frequency comparison has been achieved using well-established optical frequency transfer techniques by directly disseminating stable frequency signals [14, 15, 16, 25, 26], whereas the timescale comparison requires an exchange of a series of labeled optical pulses generated by frequency standards [11, 12, 13, 17, 18, 19, 20, 21, 22, 23, 24]. Since the optical two-way time transfer (OTWTT) method uses the reciprocity of bi-directional light travel to eliminate most of the common-mode noises and the link drift [31, 32], the measurement of optical pulses has become the key to improving the precision for time transfer.
@@
In the time synchronization of two microwave clocks, the electrical measurement has not been replaced by the optical approach because a phase coherence of independent optical frequency combs (called combs for simplicity) at both local and remote sites, which is a necessary requirement for performing stable and low-jitter interference, is challenging to establish. Specifically, microwave-to-microwave synchronization is difficult because both reference and synchronized clocks have greater phase noise than optical clocks and the pulse coherence is reduced by propagation-induced pulse impairments. It was challenging to consider not only the impact of inferior reference sources on synchronization but also the effects of fiber on long-distance optical frequency comb transmission. These factors collectively affect the precision sharing of remote microwave clocks.
To overcome the limitations mentioned above, here, we propose a tight microwave-clock time synchronization scheme based on the dual-comb-enhanced OTWTT and provide a demonstration over 205.86 km commercial fiber. We achieved a 6.23-fs residual time deviation between synchronized timescales at 1 s and an instability below at 10,000 s due to a series of breakthroughs, including high-precision locking, LOS, dispersion compensation, robust timing discrimination, and signal processing. The intensity-encoded comb pulses were measured using a Kalman filter–assisted LOS, where the centroids of reference and target interferograms at each site served as the start and stop times of the time interval, to reduce the impact of the intensity noise and pulse distortion on the peak measurements. We used the phase time of the pulse envelope instead of the carrier phase for time synchronization [20] to remove the limit of the carrier envelope offset (CEO) frequency, since the stable carrier phase is vulnerable to fluctuations in independent sources and fiber transmission. Dispersion compensation, fully automatic polarization control, and real-time feedback are essential to guarantee the mutual optical coherence.
In contrast to the majority of reported works [12, 13, 22, 19], our research specifically addressed the long-haul time synchronization scenario by transmitting optical frequency combs through optical fibers, under the conditions of utilizing microwave clocks with inferior noise characteristics. Although our approach for time synchronization using microwave clocks as reference sources achieves slightly lower precision compared to time synchronization schemes using optical clocks, we believe that the improvement in microwave clock synchronization capabilities has greatly expanded the application range of the entire high-precision clock networks.
@@
II Concept
The main task of this work was to achieve synchronization of optical pulses carrying time information from two microwave clocks, one located at the master site (site A) and the other at the synchronized site (site B). Site A and site B are adjacent to each other and located at one laboratory with independent rubidium clocks to easily quantify the residual transfer noise. Note that the microwave clock in this paper refers to both the microwave oscillator (a rubidium clock) and the optical “clock” comb that was being synchronized. The overview of the microwave-clock-based time synchronization via fiber is given below, and the specific experimental layout will be introduced in the next section.
@@ II Concept
(1) Creating the local timescales. At each site, a local timescale is created by phase-locking the repetition frequency of the comb to the local microwave clock, which served as the clock comb. Then, a comb, which is also locked to the local clock, output optical pulses with a repetition frequency offset by relative to the clock comb, which serves as the local oscillator (LO) comb to sample clock combs for performing LOS.
(2) Generating band-limited pulses. The international-telecommunication-union-standard (ITU) dense wavelength division multiplexing (DWDM) module is applied to the output of each comb to generate band-limited pulses with different spectra. The LO comb is filtered to generate two outputs with a bandwidth of , which are centered at and . The clock comb is then filtered to generate an output centered at (for site A) or (for site B).
@@
(3) Elimination of the ambiguity. Since directly using the LOS for synchronization would introduce the ambiguity of , which would cause cycle slips, we provide an unambiguous synchronization method by encoding timemarkers on the optical pulse using Mach–Zehnder modulator (MZM) (Methods). For each site, the output after the intensity modulator is separated into two parts. One part is sent to the link for transmission; the other aims to generate the local reference interferogram.
(4) Dispersion compensation. Link disersion is compensated by using dispersion compensating fibers (DCFs) at the repeater station and both receiver sites. Especially, a small amount of DCF at each receiver site is used to compensate for the residual dispersion in a fine-tuned manner. For more precise dispersion compensation, techniques such as liquid crystal modulation could be used [40].
(5) Dual comb–based LOS measurement. At each site, two cross-correlations are generated. The reference interferogram is generated by the cross-correlation between the LO comb and the local clock comb. The target interferogram is generated by the cross-correlation between the LO comb and the incoming clock comb from the remote site, which correspond to the “time-stretched” start and stop signals of the measuring time intervals, respectively. The equivalent timescale stretching parameter is obtained as . Then, an analog to-digital converter (ADC) is implemented to digitize the train of interferograms. The data is subsequently transmitted to a field programmable gate array (FPGA) controller and synchronized with the local clock for data processing.
(6) Timing discrimination. The centroid of the interferogram is used to discriminate the temporal position of the optical pulses, defined by
for the -th labeled interferogram. Here, is the envelope extracted by the Hilbert transformation, since it is not sensitive to the pulse shape and can smooth out the intensity noise superimposed on optical pulses. Then, the time difference is obtained by comparing the centroid position of two interferograms.
(7) Kalman filter–assisted LOS measurement. The Kalman filter is then utilized to suppress measurement noise.
(8) In-loop time offset calculation. The real-time timing discrimination of both sites and the calculation of the in-loop time offset are performed in the two FPGA controllers. First, at both sites, the respective time differences under a laboratory timescale and are obtained by locally comparing the centroids of the labeled reference and target interferograms using the time-difference calculation algorithm. The effective (actual) time differences at both sites are and . Then, the local generated time difference at site A is transmitted to site B using a modulated laser over the DWDM communication link. Combined with the local calculated time information , the in-loop time offset is calculated in the FPGA controller of site B.
(9) Remote clock synchronization. The time offset calculation, which acts as the error signal, is fed into the digital-feedback system that contained a proportional integral loop filter and a direct digital synthesizer (DDS). The synchronization is achieved by adjusting the repetition frequency of the clock comb at site B.
(10) Out-of-loop verification. To verify the time offset of the clocks after synchronization, the out-of-loop measurement of the time offset , which is independent from the in-loop calculated time offset , is performed. An output on both combs at each site is filtered to generate the out-of-loop reference interferogram that is independent from the in-loop signals to verify the synchronization results.
III Experimental layout
The experiment was performed in Wuhan commercial link. As shown in Fig. 1 (a), two sites A and B are adjacent to each other and located at Wuhan laboratory. Site A and site B were connected through two geographically folded optical fibers linked in Liuxinji (round-trip node), and the total link length was 205.86 km (the total link loss was 57.18 dB). The long-haul fiber transmission necessitated the installation of two optical repeater stations at Wuhan North, each of which contained an erbium-doped fiber amplifier (EDFA) and a DCF along the terrestrial link to recover the signal impairments caused by loss and dispersion. OTWTT was achieved by using the DWDM technology over the shared fiber link.
At each site, we used a dual comb–based LOS measurement to replace the TIC in the measurement of pulses generated by microwave clocks. As shown in Fig. 1 (b), at each site, we used two combs with the repetition frequency of 100 MHz and 100.001 MHz for performing LOS. The principle of LOS is shown in Fig. 1 (c). Phase locking was achieved with a locking-induced noise of 45 fs (time deviation at 1 s) through an asynchronous locking system (see Methods) to create the local timescales.
Both combs had center frequencies of 1,560 nm and a 3-dB spectrum bandwidth of approximately 15 nm. The spectral effectively covered the commonly used C-band range in fiber optic communication, which is necessary for two-way transmission. Here, we directly applied the ITU DWDM module to the output of each comb to generate band-limited pulses with different spectra. We filtered the LO comb to generate two outputs with a bandwidth of 0.8 nm (100 GHz), which were centered at 1,551.72 nm (ITU-C32) and 1,550.12 nm (ITU-C34). We filtered the clock comb to generate an output centered at ITU-C32 (for site A) or ITU-C34 (for site B).
@@ III Experimental layout
At each site, two cross-correlations were generated. The cross-correlations were generated with the equivalent sampling rate THz in the optical frequency domain and the updating rate of ms in the laboratory timescale. The equivalent timescale stretching parameter was obtained as . Then, we implemented an ADC with a 400-MHz sampling rate and a 10-bit resolution to digitize the train of interferograms, and then the data were transmitted to a FPGA controller for data processing.
@@
The fundamental task of OTWTT is to recover timing information from exchanged optical pulses. Because signal distortion and noise accumulation were caused by the long-haul fiber transmission, high-precision timing discrimination was challenging. Instead of the commonly used heterodyne-voltage-to-time scaling [12], peak location [13] or carrier phase [20] to define the arrival time in free-space scenarios, we used the centroid of the interferogram to discriminate the temporal position of the optical pulses, defined by Eq. (1).
@@
In contrast to the signal fading from turbulence in the free-space scenario [12, 19], the decrease in visibility of the target interferogram mainly attributes to the polarization disturbance in fiber, which is eliminated by the in-line electronic polarization controller. Therefore, the hold-over behavior of the Kalman filter was not considered in our experiments. In this paper, we focused on the Kalman filter to limit the amount of measurement noise [42, 43]. We assumed two zero-mean Gaussian random processes in the prior noise model of Kalman filter to represent the random-walk phase noise and random-walk frequency noise for real-time data processing [22], because these two noises are usually dominant in the microwave clock. This clock model is often applied to analyze the clock offset and Kalman-filter-based clock synchronization [42, 43, 44].
@@
IV Results
IV.1 Noise floor of the LOS measurement
Figure 2 (c) shows noise power spectral densities (PSDs) of the measurement process. The effective bandwidth of the Kalman filter was 0.01 Hz, and the PSD was reduced by approximately three orders of magnitude reduction at 0.5 Hz to 1.2 , which is close to the noise floor of the optical clock–based LOS measurement [12, 13, 22]. In addition, Figs. 2 (d) and (e) show that the Kalman filter–assisted LOS optimizes the measurement instability for a short averaging time. The shaded region shows the noise suppression below 100 s. In Fig. 2 (d), the time deviation of noise floor was optimized to 3.7 0.1 fs for 1 s of averaging time, which can support the measurement of femtosecond-level fiber changes. In Fig. 2 (e), the modified Allan deviation (MDEV) changed the slope of the noise floor from to after the Kalman filtering ( is the averaging time in seconds). Thus, the dominant white phase measurement noise was suppressed to the lower flicker phase noise region at short timescales (below 100 s), which determined the noise limitation of the time offset measurement.
IV.2 Free-running link
We then used our system to investigate the one-way free-running link. Specifically, we set both sites that shared a common master clock to test the open-loop link noise from A to B. After passing through the link, the incoming clock comb accumulated link effects, which resulted in the pulse impairments. One impairment is dispersion, which caused pulse broadening and distortion. The results show both second-order dispersion, which caused pulse broadening, and third-order dispersion, which resulted in pulse asymmetry and distortion. Additionally, the link noise and amplified spontaneous emission (ASE) noise introduced by EDFAs inevitably introduced intensity fluctuations in the optical pulses. Both dispersion-induced noise and ASE-induced intensity fluctuations decreased the pulse coherence, which ultimately affected the measurement jitter of the optical cross-correlation.
@@ IV.2 Free-running link
Imperfect dispersion and nonlinearities led to waveform distortion, which mainly affected the estimation of time offset in two aspects. First, there was a fixed systematic offset relative to the ideal position of centroid, which guided us to strictly calibrate the offset after the imperfect compensation. Second, these could lead to insufficient concentration of signal energy near the centroid, which could result in larger fluctuations when calculating the envelope and centroid using limited data. Therefore, it is important to carefully select both the amount of data and the data region used for calculating the centroid when the compensation was not perfect. We will leave the investigation of the impact of imperfect dispersion compensation and nonlinearities on system performance for future work.
IV.3 Synchronized link
After fully characterizing the 205.86-km fiber link, by performing an OTWTT to cancel the link’s influence and closing the synchronized feedback loop, we obtained the performance of the synchronized link, which was continuously acquired for over 24 hours. The synchronized link in our work, distinctly from the active stabilization link mentioned in Ref. [49], refers to the pulse-synchronization link achieved after eliminating time deviations through feedback systems. As shown in Fig. 1 (d), the time offset calculation, which acted as the error signal, was fed into the digital feedback system that contained a proportional integral loop filter and a direct digital synthesizer (DDS). It also had a 30-Hz synchronization bandwidth to force the minimized residual timing jitter of the clock output to synchronize two microwave clocks. Note that the synchronization was actually achieved between the repetition frequency of the remote clock comb and the local clock comb, with each piece of time information carried from its local clock, respectively.
@@ IV.3 Synchronized link
The residual MDEV between remote parties show the performance of the time-frequency transfer in Fig. 3 (b). The MDEV we provide is not the stability of a single microwave clock, but two clocks’ relative stability after synchronization. The relative stability between two rubidium clocks is at 1 s and reaches the floor of at 2,000 s without synchronization. The residual MDEV began at at 1 s and reached below at 10,000 s. The residual MDEV was below the state-of-the-art microwave clocks [45, 46, 47, 48] after 1 s of averaging time and more than two orders of magnitude better after the long averaging time. The slope was below 1,000 s, which indicates that the dominant noise at short timescales is flicker phase noise, and the residual MDEV follows power laws after 1,000 s averaging time. All of the results are consistent with the time deviation.
We also compared the stability result of in our experiment with the free-space scenarios that synchronized the optical clocks. The stability was approximately one order of magnitude less than that of Ref. [13], given by a few . This lower starting point was mainly due to the higher noise floor of the LOS measurement and ASE noise of the in-line EDFAs, which relatively weakened the coherence of optical pulses and caused a larger timing jitter in the microwave-clock-synchronization scenarios than in the optical oscillator–based scenarios. Moreover, compare to the result of in Ref. [12], the trend of comes from the flicker phase noise caused by the technical noise of devices, so the delay-unsuppressed white-phase noise especially with low frequency [50, 51] could not be detected. Thus, we should optimize the microwave clock–based LOS measurement. Nevertheless, this result, which is close to the optical clock time synchronization, greatly closes the gap of microwave-microwave time synchronization in the entire time-frequency network.
Recently, similarly to what we did, H. J. Kang et al. [52] utilized the filtering comb with WDMs to transmit multiple optical carriers in free space scenarios, which also provides a new possibility for comb-based time-frequency transfer.
IV.4 Probing fiber delay non-reciprocity
Similar to the probe of non-reciprocity from multi-path effects [12] or motion platforms [19] in free space, many impacts can affect the non-reciprocity in fiber-based OTWTT, which affects the true clock-offset estimation results. A particular example of the recently extensively concerned potential asymmetric delay attacks [37, 53, 39]. The improvement of the clock-offset measurement enables us to sense the subtle delay non-reciprocity of the fiber link and improves the security of long-haul fiber transmission.